A certain charge is to be divided into two parts (Q-q) and q. What is the relation Q to q if the two parts, placed a given distance apart, are to have a maximun Coulumb repulsion?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Force = K(Q-q)q/r^w
Now Numerator = K(Q-q)q should be maximum Hence,
q ahould be minium.
Rherfore,
dN/dq = k[Q-q-q]  = 0
Q – 2q = 0
Q = 2q is tthe relation.